OpenStudy (anonymous):

While simplifying some math work, Peter wrote on his paper that x^3 • x^3 • x^3 • x^3 equaled x^3(+ 3 + 3 +3) . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x^3 + x^3 + x^3 + x^3?

OpenStudy (anonymous):

Now, I know x^3 * x^3 * x^3 * x^3 = x^12 and x^3 + x^3 + x^3 + x^3 = 4x^3 But what about x^3 +^3 +^3 +^3 and how can a radical be simplified further?

OpenStudy (anonymous):

it can't be simplifies it will be just 4x^3

OpenStudy (anonymous):

I'm not worried about the x^3 + x^3 + x^3 + x^3 since I put the answer up there. I am looking for x^3 +^3 +^3 +^3

OpenStudy (anonymous):

@UditKulka